OpenStudy (anonymous):

Write in Standard form of the Quadratic Equation: i(t)=t^2-8t

OpenStudy (anonymous):

i(t)=(t-4)^2-16

OpenStudy (anonymous):

complete the perfect square now... when i(t)=84

OpenStudy (anonymous):

(t-4)^2=84+16 (t-4)^2=100 is that enough?

OpenStudy (radar):

t^2 - 8t - i(t) = 0 (might be)